OpenStudy (anonymous):

how many cm in 3.25 km?

OpenStudy (katrinakaif):

325000 cm

OpenStudy (anonymous):

i know the answer, i need an explanation

OpenStudy (saifoo.khan):

Use the ratio method.

OpenStudy (lukecrayonz):

Use a conversion chart. It can easily be remembered by saying 'King Henry Died By Drinking Chocolate Milk'. King = Kilo, Henry = Hecta, D = Deca, By = Base (Grams, meter, liter), Drinking = Deci, Chocolate = Centi,, Milk = Milli. If you have ten grams, then you have 100 decigrams. and 1000 centigrams, and 10,000 milligrams. You'd also have 1 Decagram, 0.1 hectagrams, and 0.01 kilograms.

OpenStudy (katrinakaif):

Do you know the "King henry died of drinking chocolate milk?"

OpenStudy (anonymous):

i know the answer, i need an explanation

OpenStudy (anonymous):

on my test i willl be given the chart... i just don't know how to convert it

OpenStudy (katrinakaif):

You will have to put the number in the "b" place..and move the decimal to where you want to reach it..as in lets say "K" (kilo meter) then move the decimal place as how many times you need to jump from "B" to "K"

OpenStudy (anonymous):

I'd go with the ratio method, like Saifoo says:\[\#cm = \#km \times \frac{1000m}{1km} \times \frac{100cm}{1m}\]

OpenStudy (katrinakaif):

I am expalining in baby words =/

OpenStudy (anonymous):

Notice how the units of km, and m cancel out,

OpenStudy (lukecrayonz):

When you have 1 kilometer, and you have to find how many meters, you move the decimal point to the RIGHT, as you are finding MORE, and are going down the scale. If you are finding how many kilometers are in one meter, you move the decimal LEFT as you are finding less. So, 1 meter = .001 Kilometer, 1 kilometer = 1,000 meters.

OpenStudy (katrinakaif):

-Whatever method you feel comfortable using. Its always easier to remember with a sentence that goes along with it.

OpenStudy (anonymous):

i know the answer, i need an explanation

OpenStudy (anonymous):

i'm sorryy i still don't understand... i'm so bad at this. i'm going to have to convert ml to L, cm to miles, cm to km... etc i just need a simple explanation on how i can do these

OpenStudy (lukecrayonz):

Okay. I'll draw something.

OpenStudy (lukecrayonz):